Find the slope of the line, if it is defined.
Through (-3, -9) and (5, 4)

Answer by nerdybill(7384)

Find the slope of the line, if it is defined.
Through (-3, -9) and (5, 4)
.
m (slope) = (y2-y1)/(x2-x1)
m (slope) = (4-(-9))/(5-(-3))
m (slope) = (4+9)/(5+3)
m (slope) = (13)/(8)
m (slope) = 13/8